25 mL of a 0.1111 M NaCl solution is pipetted into a clean dry 250 mL beaker. To this 25 mL of deionized water is also pipetted.
The concentration of NaCl in the 250 mL beaker is 0.1111 M. True or false?
1
Expert's answer
2019-09-19T03:59:23-0400

No. It is not true.

Cm=n/V. n=Cm x V = 0.1111 x 0.025 =0.0027775 moles.

Cm = 0.0027775 / 0.05 = 0.05555 M